# Reference-Based Post-OCR Processing with LLM for Precise Diacritic Text in Historical Document Recognition

### v1

@@ 1 Introduction
Although optical character recognition (OCR) for pervasive languages such as English, utilizing deep learning methodologies, has attained a great performance for digitizing book-form documents, the research community focus has thus transitioned to new challenges. Conversely, for languages replete with diacritics, such as Vietnamese (67 diacritic letters), Slavic (25 letters), Czech (19 letters) (Stankevičius et al. 2022), perfection remains unrealized. The problem becomes challenging when these tiny strokes of diacritics faded in historical documents like books or newspapers, and several background problems emerge such as ripped, taped, obscure words, ink smeared, watermarked, sealed by third parties, termite bitten, and bleed-through.
@@
Several projects (Gooding, Terras, and Warwick 2013) digitize these aged books by capturing page images to preserve them electronically. The next step is to transcribe images for text-based storage and later text mining, which is also a time-consuming process. Many earlier industrial tools and modern AI models were developed for this purpose. However, the precision and quality of the text generated automatically remain inadequate for uncommon languages. So, proofreading and manual modifications are necessary following these processes. To alleviate this issue, researchers leveraged the well-known task of spelling check from office text editing to correct post-OCR texts similarly. Some are based on statistical approaches to offer a list of candidates for replacing words, while recent methods offer full-text (e.g., paragraph) transformation based on the context and distribution the model was trained beforehand (Hládek, Staš, and Pleva 2020).
@@
Nevertheless, the kind of self-sufficiency spell correction model has some limitations. Let’s take the letter ‘a’ in Vietnamese, it has 17 variants with diacritics, the same thing happened with their ‘o’ letters. So, “con ga” could have $18\times 18$ combinations (assuming consonants are correct) and 10 of them are meaningful. Therefore, one limitation of the standalone spell correction model is that it heavily depends on OCR capability because the more errors are made (vowels, consonants, diacritics), the more combinations and candidates are created (Hládek, Staš, and Pleva 2020). This leads to the less chance of generating correct text lastly. The second problem especially comes with seq2seq style models, which were trained on recent corpus. Those trained sets have a different distribution (modern contents) compared with the distribution of classical contents where archaic, obsolete vocabulary and old-fashioned grammar exist. Some tried to directly train from suitable distribution but with the cost of annotation data.
@@
By chance, we discovered that, before recent advancements in AI, from the 1990s, readers and preservation enthusiasts made efforts to digitize aged books to electronic form. Ebook makers usually focus on the main content of the book as it is the final form fitting on ebook readers which provides flexibility to the screen size where pagination is less important. Thus, immediate processed records (i.e. page transcriptions) were either discarded or stored with less diligence on local storage and lost over time due to hardware failures and the absence of widely adopted personal cloud storage until around 2012.
@@
This situation occurs within a Vietnamese forum [^1] https://tve-4u.org/ where book enthusiasts collaborate to transcribe, proofread, and produce ebook files, without page-separated transcriptions. It is noteworthy that Vietnamese classical document records were published during a complex transitional period in history, leading to non-standardized grammar and orthography compared to contemporary norms. Consequently, similar yet distinct words may be used synonymously by different authors from various regions, besides, texts contain archaic, obsolete lexis. Fortunately, participants in the preservation attempt to preserve the original orthography of the texts. These ebook files are hypothesized to be valuable reference sources for document digitization and later downstream tasks. Similar sites also exist in other languages such as French e.g., Gutenberg [^2] https://www.gutenberg.org/ebooks/bookshelf/313
@@
To achieve decent accuracy, the supervised OCR method is preferred; however, we cannot directly leverage the content-focus ebooks for training, as they are not page-to-page (page2page) mapping. To solve the issue, we propose a method with the goal of using these ebooks to eventually obtain a reliable page2page label. In general, we first reduce noises by a few heuristic filters. Next, consider content-focus ebooks as reference sources, imperfect OCR-generated text then be corrected by a large language model (LLM) based on semantic and visual similarity between them and the reference. The final pseudo label is returned after short-length, dangling texts are adjusted by LLM-based spelling correction with same-page context. Details of the full pipeline are described in Section 4.
@@
To verify our hypothesis, we gathered a collection of rare, aged books published from 1850 onward, along with corresponding ebook files. This effort also served to generate a high-precision benchmark dataset for Vietnamese, as a public dense-text dataset of books with historical characteristics is still not available. Based on an evaluation of a subset of the dataset, our post-processing yields better quality labels with an evaluation rate of $8.72/10$, compared to only $7.03/10$ for the state-of-the-art transformer-based spell correction approach. We also trained 2 baselines using our final pseudo labels to compare with contemporary models like GPT-4o, Google Document AI, and Tesseract. Results at this point showed that GPT-4o achieved the top metrics, while our baseline outperformed known open-source solutions.
@@
To summarize, in this work, we proposed a novel method to utilize available ebook resources as a reference base to correct post-OCR text. The method employs LLMs to generate fine-grained pseudo ground truth without additional data annotation costs while maintaining the historical characteristics of classical documents. The methodology is applicable not only to Vietnamese but also to other diacritical languages, creating a basis to generate a quality page2page dataset to train a better open-source OCR model for underserved communities. A second main contribution is, that we introduce the first large Vietnamese publicly available dataset named as VieBookRead for classical books from the 19th century, which possess unique characteristics that are uncommon in contemporary works to close the gap in existing datasets and facilitate later research.
@@ 2.2 Spelling, Diacritics, and Typos Correction
Regarding the post-processing after the OCR and document analysis step, spelling (typos) correction or diacritics restoration techniques are sometimes leveraged to improve text recognition results in many diacritic languages. (Ruch, Baud, and Geissbühler 2003) proposed lexical disambiguation and named-entity recognition to correct spelling in the patient record in French; while (Atserias et al. 2012) used lexical co-occurrence to decide which word variants for a given context to perform spell check in Spanish. Recently, (Stankevičius et al. 2022) utilized a ByT5 Transformer to perform diacritics restoration on benchmark datasets of 12 languages such as Croatian, Latvian with high word-level accuracy.
@@ 3.3 Comparison with Other Datasets
Our dataset in red data points shows a number of areas of coverage beyond what has been achieved in previous research. Additionally, Table 1 compares some aspects of datasets: our VieBookRead has a comparable large size of 25,585 pages of book images written mainly in Vietnamese but also in other 3 languages occasionally. The features pose a uniqueness compared with other datasets which are scene texts, Nom scripts, and book covers already covered in the previous works. Some samples are included in the Appendix.
@@ 4.5 Localize Reference Text Region for Generated Paragraphs
Before clarifying this procedure, it is noteworthy to acknowledge that scanned pages occasionally contain stranded content (dangling text) where ebook creators have either neglected to transcribe the text or intentionally skipped it. This is because of various reasons such as explicit content, sensitivities, or censorship demands. Consequently, certain text segments extracted via OCR may not correspond to any content within the associated ebook.
@@ 4.6 Text Correction using LLM with Localize Reference Text Region
Given an html content as below:
{reference_html}
Providing a text which may need correction as it is generated from an imperfect OCR tool.
Use html content as reference source to correct below input text if needed:
<input-text>{paragraph}</input-text>
Follow orthography, diacritic and punctuation of the reference, only correct words existed within input text!
(other instructions for less hallucination and well-formed answer)
@@ 4.7 Minor Spelling Correction Given Context
<input-text>{text}</input-text>
Provide adjacent content nearby the input:
<content>{nearby}</content>
Correct spelling only if needed (just minimal change!)
Return the original if it is already fine (or it is random nonsense)
(other instructions for well-formed answer)
Prompt 2: Minor Spell Correction by LLM
@@ 5 Experiment
In implementation, we chose OpenAI GPT-4o mini (2024-07-18) as the LLM for all the relevant tasks of our pipeline to build the dataset, as it is the most cost-efficient model in the market having understanding in non-English languages such as Vietnamese, French. We used their batch API which is non-real-time at reduced price with a 0 temperature for more consistent answers and minimal cost for the project. Other details of parameters used in our pipeline are described in the Appendix.
5.1 Sampled Subset Evaluation
As the large number of samples in our dataset with a limited resource, we are only able to evaluate the quality of our final dataset through a subset. We conducted a stratified random sampling each of 123 books with 1025 samples for evaluation. We distributed to evaluators who received and understood the guidelines on how to evaluate. Evaluators who know Vietnamese were asked to see the page, and then grade transcriptions from 1 to 10 from no-match-at-all to perfect respectively. A perfect score (10) is given for no error or less than 2 subtle errors at word level; later 1 penalty is applied for every 2 errors from grade 9 to 6; then 1 penalty for every 3 errors from grade 5 to 2. Due to participant availability, 450 samples were ultimately evaluated.
We averaged the score given by each evaluator and compared ours with (Tran et al. 2021) transformer-based approach (publicly available and recent). Our approach scored a mean of 8.72 with a median of 8.8, SD$=0.63$ while (Tran et al. 2021) obtained a mean of 7.03 with a median of 7.46, SD$=0.94$. This suggests our method yields a better pseudo label compared to the transformer-based spelling correction. Note that (Tran et al. 2021) approach was input by our noise-suppressed texts for a fair comparison.
@@ 5.2 Baseline Models
To support the Donut baseline, we built a synthetic set by sampling blank pages with various noise backgrounds. Then, we used a GPT-2 (Radford et al. 2019) based model fine-tuned [^3] https://huggingface.co/NlpHUST/gpt2-vietnamese on a Vietnamese corpus to generate long texts from Vietnamese prompts in various fields. These texts were synthesized to the blank pages with varied layouts and styles (e.g. font, positions) to create 100,000 samples.
5.3 Evaluation
We use common OCR task metrics listed below to evaluate:
  • Edit distance we report the normalized edit distance.
  • F1 score is the harmonic mean of precision and recall.
  • BLEU (Papineni et al. 2002) measures machine translation quality but it can also be used for evaluating OCR.
  • CER Character error rate indicates the percentage of characters that were inaccurately predicted.
  • WER Word error rate indicates the percentage of words that were inaccurately predicted.

### v3

@@ 1 Introduction
Optical character recognition (OCR) utilizing deep learning methodologies for pervasive languages, such as English, has attained great performance for digitizing book-form documents, and the research community focus has thus transitioned to new challenges. Conversely, for languages replete with diacritics, such as Vietnamese (67 diacritic letters), Slavic (25 letters), Czech (19 letters) (Stankevičius et al. 2022), hopes of similar performance remain unrealized. The problem becomes challenging when these tiny strokes of diacritics faded in historical documents like books or newspapers, and several background problems emerge such as ripped, taped or obscured words, smeared inks, watermarks, third party seals, termite bites and bleed-through. Digitization and preservation of historical texts create substantial social impact, especially as social science researchers and heritage offices struggle to reliably collect and retrieve the data needed for their work, research, and education activities.
@@
Several projects (Gooding, Terras, and Warwick 2013) digitize these aged books by capturing page images to preserve them electronically. The next step is to transcribe images for text-based storage and later text mining, which is also a time-consuming process. Many earlier industrial tools and modern AI models were developed for this purpose. However, the precision and quality of the text generated automatically remain inadequate for uncommon languages. So, proofreading and manual modifications are necessary following these processes. To alleviate this issue, researchers leveraged the well-known task of spelling check from office text editing to correct post-OCR texts similarly. Some are based on statistical approaches to offer a list of candidates for replacing words, while recent methods offer full-text (e.g., paragraph) transformation based on the context and distribution the model was trained on beforehand (Hládek, Staš, and Pleva 2020).
@@
Nevertheless, the kind of self-sufficient spell correction model has some limitations. Let’s take the letter ‘a’ in Vietnamese, it has 17 variants with diacritics, the same is true for their ‘o’. So, “con ga” could have $18\times 18$ combinations (assuming consonants are correct) and 10 of them are meaningful. Therefore, one limitation of the standalone spell correction model is that it heavily depends on OCR capability because the more errors are made (vowels, consonants, diacritics), the more combinations and candidates are created (Hládek, Staš, and Pleva 2020). This leads to a lesser chance of generating correct text. The second problem especially comes with seq2seq style models, which were trained on recent corpus. Those trained sets have a different distribution (modern contents) compared with the distribution of classical contents where archaic, obsolete vocabulary and old-fashioned grammar exist. Some tried to directly train from suitable distribution but with the cost of annotation data.
@@
By chance, we discovered that, before recent advancements in AI, from the 1990s, readers and preservation enthusiasts made efforts to digitize aged books in electronic form. Ebook makers usually focus on the main content of the book as it is the final form fitting on ebook readers which provides flexibility to the screen size where pagination is less important. Thus, intermediate processed records (i.e. page transcriptions) were either discarded or stored with less diligence on local storage and lost over time due to hardware failures and the absence of widely adopted personal cloud storage until around 2012.
@@
This situation occurs within a Vietnamese forum [^1] https://tve-4u.org/ where book enthusiasts collaborate to transcribe, proofread, and produce ebook files, without page-separated transcriptions. It is noteworthy that Vietnamese classical document records were published during a complex transitional period in history, leading to non-standardized grammar and orthography compared to contemporary norms. Consequently, similar yet distinct words may be used synonymously by different authors from various regions, besides, texts contain archaic, obsolete lexis. Fortunately, participants in the preservation attempt to preserve the original orthography of the texts. These ebook files are hypothesized to be valuable reference sources for document digitization and later downstream tasks. Similar sites also exist in other languages such as French e.g., project Gutenberg [^2] https://www.gutenberg.org/ebooks/bookshelf/313.
@@
To achieve decent accuracy, a supervised OCR method is preferred; however, we cannot directly leverage the content-focus ebooks for training, as they are not page-to-page (page2page) mapping. To solve the issue, we propose a method with the goal of using these ebooks to eventually obtain a reliable page2page label. In general, we first reduce noises by a few heuristic filters. Next, consider content-focus ebooks as reference sources, imperfect OCR-generated text can then be corrected by a large language model (LLM) based on semantic and visual similarity between them and the reference. The final pseudo label is returned after short-length, dangling texts are adjusted by LLM-based spelling correction with same-page context. Details of the full pipeline are described in Section 4.
@@
To verify our hypothesis, we gathered a collection of rare, aged books published from 1850 onward, along with corresponding ebooks. This effort also served to generate a high-precision benchmark dataset for Vietnamese, as a public dense-text dataset of books with historical characteristics is still not available. Based on evaluations of the dataset, our post-processing yields better quality labels with a human-based evaluation rate of $8.72/10$ (and 86.66% semantic better), compared to only $7.03/10$ (and 6.81% semantic better) of the state-of-the-art spell correction approach. We also trained 2 baselines using our final pseudo labels to compare with contemporary models like GPT-4o, Google Document AI, and Tesseract. Results at this point showed that GPT-4o achieved the top metrics, while our baseline outperformed known open-source solutions.
@@
To summarize, we proposed a novel method to utilize available ebook resources as a reference base to correct post-OCR text. The method employs LLMs to generate fine-grained pseudo ground truth without additional data annotation costs while maintaining the historical characteristics of classical documents. The methodology is applicable not only to Vietnamese but also to other diacritical languages, creating a basis to generate a quality page2page dataset to train a better open-source OCR model for underserved communities. A second main contribution is, that we introduce a large Vietnamese publicly available dataset named as VieBookRead for classical books from the 19th century, which possess unique characteristics that are uncommon in contemporary works to close the gap in existing datasets and facilitate later research.
@@ 2.2 Spelling, Diacritics, and Typos Correction
Regarding the post-processing after the OCR and document analysis step, spelling (typos) correction or diacritics restoration techniques are sometimes leveraged to improve text recognition results in many diacritic languages. (Ruch, Baud, and Geissbühler 2003) proposed lexical disambiguation and named-entity recognition to correct spelling in the patient record in French; while (Atserias et al. 2012) used lexical co-occurrence to decide which word variants for a given context to perform spell check in Spanish. Recently, (Stankevičius et al. 2022) utilized a ByT5 to perform diacritics restoration on benchmark datasets of 12 languages such as Croatian, Latvian with high word-level accuracy.
@@ 3.3 Comparison with Other Datasets
Our dataset in red points shows a number of areas of coverage beyond what has been achieved in previous research. Additionally, Table 1 compares some aspects of datasets: our VieBookRead has a comparable large size of 25,585 pages of book images written mainly in Vietnamese but also in other 3 languages occasionally. The features pose a uniqueness compared with other datasets which are scene texts, Nom scripts, and book covers already covered in the previous works. Few samples are included in the Appendix.
@@ 4.5 Localize Reference Text Region for Generated Paragraphs
Before clarifying this procedure, it is noteworthy to know that scanned pages occasionally contain stranded content (dangling text) where ebook creators have either neglected to transcribe the text or intentionally skipped it. This is because of various reasons such as explicit content, sensitivities, or censorship demands. Consequently, certain text segments extracted via OCR may not correspond to any content within the associated ebook.
@@ 4.6 Text Correction using LLM with Localize Reference Text Region
Given an html content as below: {reference_html}
Providing a text which may need correction as it is generated from an imperfect OCR tool. Use html content as reference source to correct below input text if needed: <<<input-text>>>{paragraph}<<</input-text>>>
Follow orthography, diacritic and punctuation of the reference, only correct words existed within input text! (other instructions for less hallucinated and well-formed answer)
@@ 4.7 Minor Spelling Correction Given Context
<<<input-text>>>{text}<<</input-text>>>
Provide adjacent content nearby the input: <<<content>>>{nearby}<<</content>>>
Correct spelling only if needed (just minimal change!) Return the original if it is already fine (or it is random nonsense) (other instructions for well-formed answer)
@@ 5 Experiment
In implementation, we chose OpenAI GPT-4o mini (2024-07-18) as the LLM for all the relevant tasks of our pipeline to build the dataset, as it is the most cost-efficient model having understanding in non-English languages such as Vietnamese, French. We used their batch API at reduced price with a 0 temperature for more consistent answers and minimal cost for the project. Other details of parameters used in our pipeline are described in the Appendix.
5.1 Dataset Quality Evaluation
To evaluate our Post-OCR processing, we set up 2 assessments: a human-critic evaluation on a subset, and an automatic evaluation for the whole dataset.
Due to the large volume, we manually evaluated only a stratified random subset. Specifically, we sampled 1025 instances from 123 books, which were then reviewed by human evaluators who had received and understood properly the guidelines. Vietnamese evaluators were asked to see the page, and grade transcriptions from 1 to 10 from no-match to perfect respectively. A perfect score (10) requires up to 1 subtle word-level error. From scores 9 to 6, every 2 errors reduce the score by 1, and from scores 5 to 2, every 3 errors reduce it by 1. Due to availability, only 450 samples were ultimately evaluated. We averaged the scores given by evaluators and compared ours with (Tran et al. 2021) approach which publicly available. Our approach scored a mean of 8.72 with a median of 8.8, SD$=0.63$ while (Tran et al. 2021) obtained a mean of 7.03 with a median of 7.46, SD$=0.94$.
@@ 5.1 Dataset Quality Evaluation
An automatic way was also proposed to evaluate relatively the whole set: we prompt GPT-4o to compare carefully 2 versions of texts by ours and (Tran et al. 2021) spell-correction with explanation; with criteria is that the better means the transcript has more sense (clearer semantics), regardless the language dialect. The result is 86.66% ours is better, only 6.81% (Tran et al. 2021) result is better while 6.5% are equal. These assessments suggest our method yields a better pseudo label compared to the transformer-based spelling correction. Note that (Tran et al. 2021) method was input by our noise-suppressed texts for a fair comparison.
@@ 5.2 Baseline Models
To support the Donut baseline, a synthetic set is created by sampling blank pages with various noise backgrounds. We used a GPT-2 (Radford et al. 2019) based model fine-tuned on a Vietnamese corpus to generate long texts in various topics. These texts were synthesized to the blank pages with varied layouts and styles to create 100,000 samples.
5.3 OCR Evaluation
We use common OCR task metrics to evaluate: Norm Edit Distance (normalized character-level discrepancies); F1 score (harmonic mean of precision and recall); BLEU (a metric from machine translation, applied to OCR (Papineni et al. 2002)); CER (Character Error Rate, percentage of incorrect characters); and WER (Word Error Rate, percentage of incorrect words).
@@ 6 Conclusion
Acknowledgments
This work was supported by the Innovative Human Resource Development for Local Intellectualization program through IITP (Institute of Information & Communications Technology Planning & Evaluation) grant (IITP-2024-RS-2020-II201489, 50%) and the IITP-ITRC (Information Technology Research Center) (IITP-2024-RS-2023-00259703, 50%) grant funded by the Korea government (MSIT).
The work was also supported by Hyundai Motor Chung Mong-Koo Global Scholarship to Thao Do (1st author) and An Vo (3rd author). We are also thankful for the API research credits from OpenAI to An Vo and Google Cloud Platform to Thao Do.